Consider the following.

10, 9, 17, 25, 27

Compute the population standard deviation of the numbers. (Round your answer to one decimal place.)

(a) Add a nonzero constant c to each of your original numbers and compute the standard deviation of this new population. (Round your answer to one decimal place.)


(b) Use the results from part (a) and inductive reasoning to state what happens to the standard deviation of a population when a nonzero constant c is added to each data item.

A)The standard deviation remains the same.

B)The standard deviation becomes c.    

C)The standard deviation decreases by c.

D)The standard deviation increases by c.

E)The standard deviation becomes 0.
Answer #1

a)

=   X / n = 88 / 5 = 17.6

= sqrt [ X2 - n 2 / n ]

= sqrt [ 1824 - 5 * 17.62 / 5 ]

= 7.4

b)

If we add constant C to each sample, the deviation remains same for new sample .

Standard deviation = 7.4

c)

The standard deviation remains the same.